# In-Context Learning Enhanced Credibility Transformer

### v1

@@ 5.1 Motivation for zero-shot prediction
Zero-shot prediction is used here to refer to the ability of a model to make accurate inferences on categories or feature levels that were entirely absent during model training. In practical insurance applications, this capability is crucial, as new categories may emerge over time – for example, a new vehicle model introduced by a manufacturer or an insurance product expanded to a previously uncovered geographical region. Traditional models often struggle with such unseen categories, thus, in practice, often one defaults to simple imputation of a mean or median value, or setting the categorical level to a “missing” indicator. In contrast, the ICL mechanism in the Credibility Transformer architecture enables the model to leverage contextual similarities from related instances, potentially generalizing to these novel cases without retraining; it is exactly this ability that we test in this section.
@@ 5.2 Set-up for zero-shot prediction
We focus on the ‘Region’ categorical feature from the French MTPL dataset and design a zero-shot experiment as follows. We create a new train-test split where certain regions are treated as completely “unseen”. Specifically, we select records comprising approximately 10% of the exposure of the entire dataset, corresponding to the regions with the lowest amounts of exposure: [’R43’, ’R21’, ’R42’, ’R94’, ’R83’, ’R74’, ’R23’, ’R22’, ’R26’, ’R25’, ’R73’]. These regions are designated as the new test set, and their labels are remapped to a new “unseen” category to simulate novelty during inference. For the remaining training data, we further set additional region levels to “unseen” to teach the original Credibility Transformer model how to handle this category, i.e., to provide a parameter estimate for the embedding level corresponding to “unseen”. For this, we select the next smallest regions in terms of exposures – [’R41’, ’R54’, ’R31’, ’R72’, ’R91’, ’R52’] – which account for roughly 20% of the exposure of the training rows. Thus, we ensure that the model encounters the “unseen” label during training, learning to rely on contextual information from similar instances. The full Credibility Transformer and ICL model pipeline are then trained on this modified dataset, and its predictive performance is assessed on the held-out test set, quantifying its zero-shot generalization to truly novel regions. Table 4 summarizes the key information of the train-test split. Interestingly, we see that the frequency in the new test set is a lower than in the training set; this is due to the new manner in which we have performed the split above.
Characteristic Training set Test set
@@
Categorical (4): Area, VehGas, VehBrand, Region
@@ 5.3 Regional Analysis
As a preliminary introduction to better understand the performance of our zero-shot approach across different regions, we present a detailed breakdown of the results by region in Table 5. This table shows the claim counts, exposure, claim rates, and Poisson deviance for each region, along with whether it was designated as “unseen” and its assignment to the training or test set. The Poisson deviance has been calculated by using the average observed rates within each region as the predictions.
@@
Table 5: Regional breakdown showing claim counts, exposure, claim rates, and Poisson deviance by region, with designation as “unseen” and assignment to training or test sets.
The regional analysis reveals interesting patterns in the distribution of risk across different geographical areas. The test regions (all designated as “unseen”) show varying claim rates ranging from 0.05 to 0.09, with corresponding Poisson deviances between 0.158 and 0.273. Notably, regions with higher claim rates tend to have higher Poisson deviances. The training regions show a similar pattern, with the “unseen” training regions generally having lower exposure volumes than the “seen” regions also used for the initial model training. These “seen” regions have a notably higher set of claims rates and Poisson deviances than the “unseen” test set.
To provide a comprehensive view of model performance across different data segments, Table 6 presents the exposure-weighted average Poisson deviances for the key portfolio segments.
Description Poisson Deviance
@@
Table 6: Weighted average Poisson deviances across different portfolio segments.
The weighted averages show that the test set achieves the lowest Poisson deviance (0.216), followed by the “unseen” training regions (0.238). Interestingly, the “seen” training regions with full regional information show the highest deviance (0.267).
The significant differences between train and test, and “seen” and “unseen” should be kept in mind when understanding the zero-shot results presented in the next section, in particular, the differences between in-sample and out-of-sample results.
@@ 5.4 Zero-shot results
To evaluate the performance of the ICL-Credibility Transformer model on unseen data, the model is first trained on the training dataset. When reporting the results of all of the modeling steps, we filter the training set to assess performance only on records labeled “unseen”. Likewise, the entire test set consists of exactly these “unseen” records, but for a different set of regions, as described above. We follow the training approach outlined in Section 4: first, we fit a base Credibility Transformer model; then, we add 2 ICL layers and freeze the decoder while training for 50 epochs; finally, in the third phase, we unfreeze all weights and train for an additional 20 epochs. All other training hyper-parameters are identical to those in Table 2. The results are displayed in Table 7.
@@
Model Param. Poisson loss Poisson loss
@@
ICL Transformer - 2 Layers (phase 2) 46,119 22.315 20.264
Fine-tuning ICL Transformer - 2 Layers (phase 3) 46,119 22.298 20.259
Table 7: Number of parameters, in-sample and out-of-sample Poisson deviance losses (units are in $10^{-2}$). Null model has been calculated to be used as a baseline
Given that this train-test split deviates from the more conventional setup shown before, we fit a null model (intercept-only) as a baseline. The base Credibility Transformer significantly outperforms the null model on both in-sample and out-of-sample Poisson deviance losses, as we would expect. Adding ICL layers to the base Credibility Transformer (with the decoder frozen) further improves the out-of-sample loss from $20.282$ to $20.264$. This gain highlights the benefits of cross-record attention enriched with contextual information, as well as the role of credibility weighting, which allows tokens to enhance their predictive power by attending to other contextually similar tokens in the batch. Finally, fine-tuning the full ICL Transformer (with the decoder unfrozen) yields an additional out-of-sample improvement to $20.259$. This indicates that the model benefits from adapting its decoder to the newly enriched context provided by the ICL to the CLS tokens.
Thus, we observe that the ICL-Credibility Transformer model generalizes fairly well to unseen levels, provided that it receives some training exposure to handling such unseen levels. In our example, we sacrifice information on a few low-exposure levels to allow the model to learn how to treat these missing levels. However, this approach does result in a loss of data. Alternative methods for including these levels might involve creating duplicates of records designated as “unseen” while retaining both the original and “unseen” labels. Care must be taken, however, to ensure that these training records do not serve as context for one another, as this could cause the model to perform worse on truly unseen test cases, i.e., such a leakage of information must be avoided.
@@ 6 Discussion and Conclusions
Looking forward beyond this initial exploration of ICL within an actuarial context, several practical issues remain. The performance of our model depends on the quality and relevance of the context batch. From a process perspective, while we employed a nearest-neighbor search in the CLS token space, scaling this retrieval process to a real world portfolios may presents a computational challenge. More importantly, a critical hurdle for practical adoption is the justification of this dynamic, context-driven approach. Regulators will demand transparency and fairness, asking why a specific set of policies was used as context for a given individual. This requires developing a way to ensure this context does not introduce biases and results in premium rates that are actuarially sound. For policyholders, the rationale must be translatable into an explanation that their rate is fine-tuned using recent, relevant, and anonymized claims experience from a peer group to ensure maximum accuracy and fairness.
Appendix. Nearest-neighbor samples and point-level narratives
Columns shown: Stage, Rank, Distance, Exposure, Area, VehPower, VehAge, DrivAge, BonusMalus, VehBrand, VehGas, Region.
Note: We show two neighbors per stage (phase1-ct (untrained CT), pre-base, pre-decorated, pre-final, post-base, post-decorated, post-final).
Point 1 (selected_index 57559)
Target: Area A, VehPower 6, VehAge 18, DrivAge 68, BonusMalus 55, Brand B1, Gas Diesel, Region R53, Exposure 1.00.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.859284 1.00 A 6 19 69 55 B1 Diesel R53
phase1-ct 2 0.970658 0.39 A 6 17 67 55 B1 Diesel R53
pre-base 1 0.354658 1.00 A 6 19 69 55 B1 Diesel R53
pre-base 2 0.358824 0.39 A 6 17 67 55 B1 Diesel R53
pre-decorated 1 0.301852 0.39 A 6 17 67 55 B1 Diesel R53
pre-decorated 2 0.333753 1.00 A 6 19 69 55 B1 Diesel R53
pre-final 1 0.452421 0.39 A 6 17 67 55 B1 Diesel R53
pre-final 2 0.680154 1.00 A 6 19 69 55 B1 Diesel R53
post-base 1 0.471765 1.00 A 6 19 69 55 B1 Diesel R53
post-base 2 0.474374 0.39 A 6 17 67 55 B1 Diesel R53
post-decorated 1 0.474778 0.39 A 6 17 67 55 B1 Diesel R53
post-decorated 2 0.528474 1.00 A 6 19 69 55 B1 Diesel R53
post-final 1 0.544454 0.39 A 6 17 67 55 B1 Diesel R53
post-final 2 0.874919 1.00 A 6 19 69 55 B1 Diesel R53
Narrative.
Already in phase1 (untrained) and in the pre–base model the nearest neighbors match the same A/R53 diesel B1 cohort with late‑60s drivers. Outcome decoration tightens proximity (best distance $\approx 0.35\to 0.30$) while preserving fuel/region/brand; the final stages broaden slightly but retain the same actuarial signature.
Point 2 (selected_index 7794)
Target: Area C, VehPower 4, VehAge 15, DrivAge 19, BonusMalus 90, Brand B14, Gas Regular, Region R24, Exposure 0.00 (0.0027).
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.915045 0.43 A 4 11 24 85 B14 Regular R24
phase1-ct 2 1.140315 0.19 A 5 13 20 90 B14 Regular R24
pre-base 1 0.576452 0.12 C 6 17 19 90 B1 Regular R53
pre-base 2 0.838548 0.52 B 7 16 20 85 B3 Regular R24
pre-decorated 1 0.403106 0.12 C 6 17 19 90 B1 Regular R53
pre-decorated 2 0.601895 0.35 C 4 11 19 85 B1 Regular R24
pre-final 1 0.646874 0.12 C 6 17 19 90 B1 Regular R53
pre-final 2 0.648549 0.25 A 5 16 20 90 B4 Regular R24
post-base 1 0.956439 0.19 A 5 13 20 90 B14 Regular R24
post-base 2 1.015568 0.52 B 7 16 20 85 B3 Regular R24
post-decorated 1 0.787472 0.19 A 5 13 20 90 B14 Regular R24
post-decorated 2 0.909297 0.52 B 7 16 20 85 B3 Regular R24
post-final 1 0.918342 0.19 A 5 13 20 90 B14 Regular R24
post-final 2 1.042231 0.52 B 7 16 20 85 B3 Regular R24
Narrative.
Phase1 already finds B14 Regulars in R24 with similar ages and high Bonus–Malus. Pre–base introduces Regular neighbors from A/C (B1/B3), and decoration tightens distances while keeping youth and Regular fuel. In the post stages, exact B14/R24 matches reappear among the closest points; final neighborhoods remain youthful, Regular and high BM.
Point 3 (selected_index 8692)
Target: Area C, VehPower 5, VehAge 1, DrivAge 60, BonusMalus 50, Brand B2, Gas Diesel, Region R24, Exposure 1.00.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.006281 0.17 C 5 1 60 50 B2 Diesel R24
phase1-ct 2 0.018002 0.01 C 5 0 61 50 B2 Diesel R24
pre-base 1 0.018674 0.01 C 5 0 61 50 B2 Diesel R24
pre-base 2 0.037920 0.17 C 5 1 60 50 B2 Diesel R24
pre-decorated 1 0.018339 0.01 C 5 0 61 50 B2 Diesel R24
pre-decorated 2 0.034097 0.17 C 5 0 62 50 B2 Diesel R24
pre-final 1 0.190356 1.00 C 5 2 64 50 B3 Diesel R24
pre-final 2 0.201400 0.43 C 6 3 59 50 B1 Diesel R24
post-base 1 0.016105 0.01 C 5 0 61 50 B2 Diesel R24
post-base 2 0.033057 0.17 C 5 1 60 50 B2 Diesel R24
post-decorated 1 0.010535 0.01 C 5 0 61 50 B2 Diesel R24
post-decorated 2 0.034913 0.17 C 5 0 62 50 B2 Diesel R24
post-final 1 0.196228 0.17 C 5 0 62 50 B2 Diesel R24
post-final 2 0.204852 1.00 C 6 1 65 50 B2 Diesel R24
Narrative.
Across stages the closest neighbors are near duplicates in C/R24 with B2 diesel and ages 59–62. Decoration further tightens the best match ($\approx 0.018\to 0.011$). Final stages admit slightly older matches (64–65, same fuel/region) but remain within the same credibility cluster.
Point 4 (selected_index 9404)
Target: Area D, VehPower 4, VehAge 0, DrivAge 57, BonusMalus 50, Brand B1, Gas Diesel, Region R24, Exposure 0.40.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.000000 0.84 D 4 0 57 50 B1 Diesel R24
phase1-ct 2 0.030982 0.94 D 4 0 58 50 B1 Diesel R24
pre-base 1 0.000000 0.84 D 4 0 57 50 B1 Diesel R24
pre-base 2 0.071298 0.94 D 4 0 58 50 B1 Diesel R24
pre-decorated 1 0.000000 0.84 D 4 0 57 50 B1 Diesel R24
pre-decorated 2 0.125568 0.08 D 4 1 59 50 B1 Diesel R24
pre-final 1 0.329011 0.49 D 4 0 60 50 B3 Diesel R24
pre-final 2 0.360727 0.94 D 4 0 58 50 B1 Diesel R24
post-base 1 0.000000 0.84 D 4 0 57 50 B1 Diesel R24
post-base 2 0.065833 0.94 D 4 0 58 50 B1 Diesel R24
post-decorated 1 0.000000 0.84 D 4 0 57 50 B1 Diesel R24
post-decorated 2 0.116503 0.94 D 4 0 58 50 B1 Diesel R24
post-final 1 0.195898 0.94 D 4 0 58 50 B1 Diesel R24
post-final 2 0.308650 0.49 D 4 0 60 50 B3 Diesel R24
Narrative.
This point has near-exact duplicates (distance $\approx 0$) in D/R24 B1 diesel, and decoration leaves that structure intact. Final stages admit B3/B2 variants and slightly different ages, but the local cohort remains the same.
Point 5 (selected_index 55076)
Target: Area C, VehPower 6, VehAge 5, DrivAge 50, BonusMalus 50, Brand B1, Gas Diesel, Region R52, Exposure 0.63.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.000000 0.36 C 6 5 50 50 B1 Diesel R52
phase1-ct 2 0.159966 1.00 C 6 2 50 50 B1 Diesel R52
pre-base 1 0.000000 0.36 C 6 5 50 50 B1 Diesel R52
pre-base 2 0.052296 1.00 C 6 2 50 50 B1 Diesel R52
pre-decorated 1 0.000000 0.36 C 6 5 50 50 B1 Diesel R52
pre-decorated 2 0.050454 1.00 C 6 2 50 50 B1 Diesel R52
pre-final 1 0.242212 0.49 C 7 4 47 50 B1 Diesel R53
pre-final 2 0.263396 1.00 C 7 5 47 50 B1 Diesel R53
post-base 1 0.000000 0.36 C 6 5 50 50 B1 Diesel R52
post-base 2 0.057980 1.00 C 6 2 50 50 B1 Diesel R52
post-decorated 1 0.000000 0.36 C 6 5 50 50 B1 Diesel R52
post-decorated 2 0.054076 1.00 C 6 2 50 50 B1 Diesel R52
post-final 1 0.207730 0.36 C 6 5 50 50 B1 Diesel R52
post-final 2 0.231302 0.97 C 5 4 47 50 B1 Diesel R53
Narrative.
A highly redundant C/R52 B1 diesel segment: phase1 through post–decorated return a literal near‑duplicate as the closest match. Final stages broaden gently into R53 with similar ages and power.
Point 6 (selected_index 54529)
Target: Area C, VehPower 5, VehAge 2, DrivAge 46, BonusMalus 50, Brand B2, Gas Diesel, Region R72, Exposure 1.00.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.095599 0.45 C 5 4 46 50 B2 Diesel R72
phase1-ct 2 0.230291 0.85 C 5 0 44 50 B2 Diesel R72
pre-base 1 0.029398 0.45 C 5 4 46 50 B2 Diesel R72
pre-base 2 0.149170 0.08 C 5 3 45 50 B1 Diesel R72
pre-decorated 1 0.022880 0.45 C 5 4 46 50 B2 Diesel R72
pre-decorated 2 0.116307 0.08 C 6 4 45 50 B2 Diesel R72
pre-final 1 0.304717 0.59 C 5 1 42 50 B5 Diesel R72
pre-final 2 0.330409 0.08 C 6 1 48 50 B1 Diesel R72
post-base 1 0.030205 0.45 C 5 4 46 50 B2 Diesel R72
post-base 2 0.143507 0.08 C 5 3 45 50 B1 Diesel R72
post-decorated 1 0.022541 0.45 C 5 4 46 50 B2 Diesel R72
post-decorated 2 0.122115 0.08 C 4 0 47 50 B3 Diesel R72
post-final 1 0.226112 1.00 C 4 7 46 50 B2 Diesel R72
post-final 2 0.271345 0.08 C 6 4 45 50 B2 Diesel R72
Narrative.
Phase1 already centres on the C/R72 B2 diesel cohort in the mid‑40s. Decoration consistently tightens distances. Final stages broaden slightly to B1/B3 diesels in the same region (and occasional adjacent regions in deeper ranks), leaving the core profile unchanged.
Point 7 (selected_index 2318)
Target: Area E, VehPower 7, VehAge 2, DrivAge 61, BonusMalus 50, Brand B6, Gas Regular, Region R93, Exposure 0.21.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.031319 1.00 E 7 2 59 50 B6 Regular R93
phase1-ct 2 0.124904 0.12 E 7 1 59 50 B6 Regular R93
pre-base 1 0.043811 1.00 E 7 2 59 50 B6 Regular R93
pre-base 2 0.050928 0.47 E 7 2 60 50 B6 Regular R93
pre-decorated 1 0.048501 1.00 E 7 2 59 50 B6 Regular R93
pre-decorated 2 0.056023 0.47 E 7 2 60 50 B6 Regular R93
pre-final 1 0.221336 0.12 E 7 1 59 50 B6 Regular R93
pre-final 2 0.317985 1.00 E 7 2 59 50 B6 Regular R93
post-base 1 0.042100 1.00 E 7 2 59 50 B6 Regular R93
post-base 2 0.056163 0.47 E 7 2 60 50 B6 Regular R93
post-decorated 1 0.054033 1.00 E 7 2 59 50 B6 Regular R93
post-decorated 2 0.071394 0.47 E 7 2 60 50 B6 Regular R93
post-final 1 0.232991 0.47 E 7 2 60 50 B6 Regular R93
post-final 2 0.273311 0.12 E 7 1 59 50 B6 Regular R93
Narrative.
A very stable E/R93 B6 Regular cluster: phase1 through post–decorated return practically the same neighbors (late‑50s/early‑60s). Final stages admit a few B6 diesel comparisons further down the list, but the area/region/fuel signature remains consistent.
Point 8 (selected_index 22962)
Target: Area E, VehPower 7, VehAge 14, DrivAge 46, BonusMalus 54, Brand B2, Gas Regular, Region R82, Exposure 1.00.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.569729 1.00 E 7 13 43 54 B1 Regular R82
phase1-ct 2 0.734658 1.00 E 5 11 45 54 B3 Regular R82
pre-base 1 0.416628 1.00 E 7 13 43 54 B1 Regular R82
pre-base 2 0.561383 0.18 E 7 12 42 54 B1 Regular R82
pre-decorated 1 0.359361 1.00 E 7 13 43 54 B1 Regular R82
pre-decorated 2 0.455073 0.18 E 7 12 42 54 B1 Regular R82
pre-final 1 0.412044 1.00 E 7 13 43 54 B1 Regular R82
pre-final 2 0.478571 0.18 E 7 12 42 54 B1 Regular R82
post-base 1 0.481606 1.00 E 7 13 43 54 B1 Regular R82
post-base 2 0.521648 0.16 E 6 13 43 55 B5 Regular R82
post-decorated 1 0.348787 1.00 E 7 13 43 54 B1 Regular R82
post-decorated 2 0.409371 0.16 E 6 13 43 55 B5 Regular R82
post-final 1 0.350916 1.00 E 7 13 43 54 B1 Regular R82
post-final 2 0.517703 0.18 E 7 12 42 54 B1 Regular R82
Narrative.
A clear E/R82 Regular cohort with ages in the low‑ to mid‑40s. Decoration reliably tightens the closest match. Post stages occasionally draw B5 Regulars (or B1 diesels farther down), but region and age alignment remain strong.
Point 9 (selected_index 63739)
Target: Area E, VehPower 8, VehAge 13, DrivAge 35, BonusMalus 72, Brand B1, Gas Regular, Region R11, Exposure 0.18.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.349907 1.00 E 8 12 34 72 B1 Regular R11
phase1-ct 2 0.487421 0.85 E 6 11 34 76 B1 Regular R11
pre-base 1 0.323057 1.00 E 8 12 34 72 B1 Regular R11
pre-base 2 0.585882 0.91 E 8 11 33 72 B1 Regular R11
pre-decorated 1 0.250448 1.00 E 8 12 34 72 B1 Regular R11
pre-decorated 2 0.378647 0.25 C 5 4 31 80 B6 Regular R31
pre-final 1 0.471002 0.08 C 7 14 32 81 B6 Diesel R72
pre-final 2 0.492369 0.08 E 5 9 28 76 B5 Regular R26
post-base 1 0.284585 1.00 E 8 12 34 72 B1 Regular R11
post-base 2 0.514235 0.91 E 8 11 33 72 B1 Regular R11
post-decorated 1 0.209932 1.00 E 8 12 34 72 B1 Regular R11
post-decorated 2 0.383713 0.91 E 8 11 33 72 B1 Regular R11
post-final 1 0.363904 1.00 E 8 12 34 72 B1 Regular R11
post-final 2 0.487113 0.08 E 7 8 33 72 B1 Regular R11
Narrative.
The dominant pattern is E/R11 Regular B1 with early‑ to mid‑30s drivers. Decoration tightens substantially (best distance $\approx 0.32\to 0.21$ pre; $\approx 0.28\to 0.21$ post). Pre–final briefly explores B6 diesel (R72) and B5 Regular (R26), but post stages re‑centre on R11 Regular B1.
Point 10 (selected_index 44827)
Target: Area C, VehPower 4, VehAge 4, DrivAge 21, BonusMalus 95, Brand B2, Gas Diesel, Region R24, Exposure 0.15.
Stage Rank Dist Exp Area VP VAge DAge BM Brand Gas Region
phase1-ct 1 0.528064 0.24 C 5 2 24 95 B2 Diesel R24
phase1-ct 2 0.584082 0.16 C 5 3 21 95 B2 Diesel R24
pre-base 1 0.159332 0.16 C 5 3 21 95 B2 Diesel R24
pre-base 2 0.322642 0.01 C 5 1 19 95 B2 Diesel R24
pre-decorated 1 0.226593 0.16 C 5 3 21 95 B2 Diesel R24
pre-decorated 2 0.376376 0.36 C 5 6 20 95 B3 Diesel R53
pre-final 1 0.553746 0.44 A 6 5 24 95 B2 Diesel R24
pre-final 2 0.575868 0.66 D 7 10 23 95 B6 Regular R24
post-base 1 0.219119 0.16 C 5 3 21 95 B2 Diesel R24
post-base 2 0.402813 0.13 C 6 9 21 95 B1 Diesel R24
post-decorated 1 0.265050 0.16 C 5 3 21 95 B2 Diesel R24
post-decorated 2 0.326072 0.44 A 6 5 24 95 B2 Diesel R24
post-final 1 0.510366 0.01 C 5 1 19 95 B2 Diesel R24
post-final 2 0.623841 0.16 C 5 3 21 95 B2 Diesel R24
Narrative.
A young, high‑BM B2 diesel in C/R24 has strong near duplicates from phase1 onwards. Decoration modestly broadens the local set (occasional B3 diesel and A/R24). Final stages keep B2 diesel neighbors in R24 with small shifts in age and power, preserving the core risk profile.

### v2

@@ 5.1 Motivation for zero-shot prediction
Zero-shot prediction is used here to refer to the ability of a model to make accurate inferences on categories or feature levels that were entirely absent during model training. In practical insurance applications, this capability can be crucial, as new categories may emerge over time – for example, a new vehicle model introduced by a manufacturer or an insurance product expanded to a previously uncovered geographical region. Traditional models often struggle with such unseen categories, thus, in practice, often one defaults to simple imputation of a mean or median value, or setting the categorical level to a “missing” indicator. In contrast, the ICL mechanism of the ICL-Credibility Transformer architecture enables the model to leverage contextual similarities from related instances, potentially generalizing to these novel cases without retraining the model; it is exactly this ability that we test in this section.
@@ 5.2 Set-up for zero-shot prediction
In the following example, we focus on the ‘Region’ categorical feature from the French MTPL dataset, and we design a zero-shot experiment as follows. We create a new train-test split where certain regions are treated as completely “unseen” (unavailable during model training). Specifically, we select records comprising approximately 10% of the exposure of the entire dataset, corresponding to the regions with the lowest amounts of exposures, these are ${\cal R}_{\rm new}=\{$’R43’, ’R21’, ’R42’, ’R94’, ’R83’, ’R74’, ’R23’, ’R22’, ’R26’, ’R25’, ’R73’}\}. These regions are designated as the new test set, and their labels are remapped to a new “unseen” category to simulate novelty during inference. For the remaining training data, we further set additional region levels to “unseen” to teach the original base Credibility Transformer how to handle this category, i.e., to provide a parameter estimate for the embedding level corresponding to “unseen”. For this, we select the next smallest regions in terms of exposures $\{$’R41’, ’R54’, ’R31’, ’R72’, ’R91’, ’R52’}\}, which account for roughly 20% of the entire exposure of the training sample. Thus, we ensure that the model encounters the “unseen” label during model training, learning to rely on contextual information from similar instances in these cases. The full ICL-Credibility Transformer model pipeline is then trained on this modified dataset, and its predictive performance is assessed on the (hold-out) test set, quantifying its zero-shot generalization to truly novel regions ${\cal R}_{\rm new}$. Table 4 summarizes the key information of the train-test split. Interestingly, we see that the frequency in the new test set is a lower than in the training set; this is due to the new manner in which we have performed the split above.
Characteristic Training set Test set ${\cal R}_{\rm new}$
@@
Categorical (4): Area, VehGas, VehBrand, Region/“unseen”
@@ 5.3 Regional analysis
As a preliminary introduction to better understand the performance of our zero-shot approach across different regions, we present a detailed breakdown of the results by region in Table 5. This table shows the claim counts, exposures, claim rates, and Poisson deviances for each region, along with whether it was designated as “unseen” and its assignment to the training or test set. The Poisson deviances have been calculated by using the average observed frequencies within each region as the predictions.
@@
Table 5: Regional breakdown: Poisson deviances by Region are computed based on the sample frequency in the corresponding Region.
The regional analysis reveals interesting patterns in the distribution of risk across the different geographical areas. The test regions (all designated as “unseen”) show varying frequency rates ranging from 0.05 to 0.09, with corresponding Poisson deviances between 0.158 and 0.273. Notably, regions with higher claim frequencies tend to have higher Poisson deviances. The training regions show a similar pattern, with the “unseen” training regions having lower exposures than the “seen” regions (by construction). These “seen” regions have a notably higher set of claims frequencies and Poisson deviances than the “unseen” test set.
Description Poisson Deviances
@@
Table 6: Weighted average Poisson deviances across the different portfolio segments.
To provide a comprehensive view of the model performance across the different data segments, Table 6 presents the exposure-weighted average Poisson deviances for the key portfolio segments. The weighted averages show that the test set achieves the lowest average Poisson deviance (0.216), followed by the “unseen” training regions (0.238). Interestingly, the “seen” training regions with full regional information show the highest average Poisson deviance (0.267). The significant differences between train and test, and “seen” and “unseen” should be kept in mind when understanding the zero-shot results presented in the next section, in particular, the differences between in-sample and out-of-sample results.
@@ 5.4 Zero-shot results
To evaluate the performance of the ICL-Credibility Transformer model on the “unseen” data (regions), the model is first trained on the training sample. When reporting the results of all of the modeling steps, we filter the training set to assess the performance only on records labeled “unseen” regions. Likewise, the entire test set consists of exactly these “unseen” records, but for a different set of regions ${\cal R}_{\rm new}$, as described above. We follow the training approach outlined in Section 4: (1) first, we fit a base Credibility Transformer; (2) then, we add 2 ICL layers and freeze the decoder while training for 50 epochs; (3) finally, in the third phase, we unfreeze all weights and train for an additional 20 epochs. All other training hyper-parameters are identical to those in Table 2. The results are displayed in Table 7.
@@
Model Param. Poisson loss Poisson loss on ${\cal R}_{\rm new}$
@@
ICL-Credibility Transformer (phase 2) 46,119 22.315 20.264
Fine-tuning ICL-Credibility Transformer (phase 3) 46,119 22.298 20.259
Table 7: Number of parameters, in-sample and out-of-sample Poisson deviance losses (units are in $10^{-2}$). Null model has been calculated to be used as a baseline using the observed frequency of the training data.
Given that this training-test split deviates from the more conventional set-up shown before, we fit a null model (intercept-only) as a baseline. The base Credibility Transformer significantly outperforms the null model on both in-sample and out-of-sample Poisson deviance losses, as we would expect. Adding ICL mechanism to the base Credibility Transformer (with the frozen decoder) further improves the out-of-sample loss from $20.282$ to $20.264$. This gain highlights the benefits of cross-record attention enriched with contextual information, as well as the role of the credibility weighting, which allows tokens to enhance their predictive power by attending to other contextually similar tokens in the context batch. Finally, fine-tuning the full ICL-Credibility Transformer end-to-end yields an additional out-of-sample improvement to $20.259$. This indicates that the model benefits from adapting its decoder to the newly enriched context provided by ICL to the CLS tokens.
Thus, we observe that the ICL-Credibility Transformer generalizes fairly well to unseen the levels, provided that it receives some training exposure to handling such unseen levels. In our example, we sacrifice information on a few low-exposure levels to allow the model to learn how to treat these missing levels. However, this approach does result in a loss of data. Alternative methods for including these levels might involve creating duplicates of records designated as “unseen” while retaining both the original and the “unseen” labels. Care must be taken, however, to ensure that these training records do not serve as context for one another, as this could cause the model to perform worse on truly unseen test cases, i.e., such a leakage of information must be avoided for successful model training.
@@ 6 Discussion and conclusions
Looking forward beyond this initial exploration of ICL within an actuarial context, several practical issues remain. The performance of our model depends on the quality and relevance of the context batch. From a process perspective, while we employed a nearest-neighbor search in the CLS token space, scaling this retrieval process to real world portfolios may present a computational challenge. More importantly, a critical hurdle for a practical adoption is the justification of this dynamic, context-driven approach. Regulators will demand transparency and fairness, asking why a specific set of policies was used as context for a given individual. This requires developing a way to ensure this context does not introduce biases and results in premium rates that are actuarially sound. For policyholders, the rationale must be translatable into an explanation that their rate is fine-tuned using recent, relevant, and anonymized claims experience from a peer group to ensure maximum accuracy and fairness.
Another potential actuarial field of applications of ICL is individual claims reserving. In that case, the context can be formed by similar individual claims that are already further developed to gain more insight how a specific individual claim may evolve. In this context, there are less regulatory concerns about fairness and explainability because this mainly affects internal actuarial processes that do not directly impact individual pricing of insurance contracts.
Declaration. We have used ChatGPT to support us in the literature overview and to improve language and readability.